how many liters each of 60% acid solution and 80% acid solution must be used to produce 60 liters of 70% acid solution?
Answer by richwmiller(17219)

x+y=60,
0.6*x+0.8*y=0.7*60
0.6*x+0.8*y=42
x=60-y
0.6*(60-y)+0.8*y=42
36-0.6y+0.8*y=42
0.2*y=6
y=30 at 80%
x=60-y
x=30 at 60%
check
0.6*30+0.8*30=0.7*60
18+24=42
42=42
ok